OpenStudy (anonymous):

simplify the expression -3x6y4 ----------- x-3y-2

OpenStudy (zzr0ck3r):

\[\frac{-3x^{6}*y^{4}}{x-3y-2}\]?

OpenStudy (anonymous):

you would have to apply one of the laws of indices in this question: \[\Large \frac{ x^a }{ x^b }~=~x ^{a-b}\]

OpenStudy (zzr0ck3r):

\[\frac{-3x^{6}y^{4}}{x^{-3}y^{-2}}\]

OpenStudy (anonymous):

@zzr0ck3r yes that is the expression, sorry. i need it simplified.

OpenStudy (zzr0ck3r):

so \[-3x^{6-(-3)}y^{4-(-2)}\]

OpenStudy (zzr0ck3r):

= \[-3x^{?}y^{?}\]

OpenStudy (anonymous):

so in this case, you have \[\Large -3x ^{6-(-3)}y ^{4-(-2)}\] \[\Large -3x^9y^6\]

OpenStudy (zzr0ck3r):

or just give them the answer:)

OpenStudy (anonymous):

thank you so much guys, i really appreciate it . mind helping me on another one? its an online course, i just need to pass.

OpenStudy (anonymous):

@kausarsalley thank you for the answer, lovely !

OpenStudy (zzr0ck3r):

close this and ask another, but I would try and understand how we got this one first. then try the next on your own, then ask if you can't get it.

OpenStudy (anonymous):

thank you (: